The total value of a country's GDP is $1000. If the country's GDP grows by 3% annually, what will be the total value of the country's in 5 years?

A.

$1209.59

B.

$1159.27

C.

$1080.29

D.

$1093.20

E.

$1293.07
Answer #1

Initial GDP of a country = $1000

Annual GDP growth rate = 3%

GDP of the country after 5 years = Initial GDP * (1+ Annual GDP growth rate)^5 = $1000 * (1+3%)^5 = $1000 * 1.03^5 = $1159.27

Ans: B. $1159.27